OpenStudy (anonymous):

what is a summation notation for the series? a. -5 +2+9+16+...+261+268 b. 500+490+480+...+20+10

Mehek (mehek14):

what do you think?

OpenStudy (anonymous):

i have no idea on how to do this

OpenStudy (jhannybean):

Iwould start with b, seems a bit easier.

OpenStudy (jhannybean):

This is what you call an arithmetic sequence, where you have a common `difference` between the numbers in a sequence. Can you spot what the difference is?

OpenStudy (anonymous):

it decreases by ten

OpenStudy (jhannybean):

I've got to head off for a bit, I'll get someone on to help you @Michele_Laino please? Thank you !

Mehek (mehek14):

yes it does decrease by 10 each time

Mehek (mehek14):

how about in A?

OpenStudy (anonymous):

by 7?

Mehek (mehek14):

increase or decrease

OpenStudy (anonymous):

increasing

Mehek (mehek14):

correct

OpenStudy (anonymous):

what's next?

OpenStudy (loser66):

For a) \(a_2= a_1+7\\a_3= a_2+7= a_1+2*7\\a_4=a_3+7 = a_1+3*7\\--------------\\a_n= a_1+(n-1)7\)

OpenStudy (loser66):

Now, calculate the last term: \(268 = a_1 +(n-1) 7=-5 +(n-1) *7\) solve for n, you have n= 40 Hence the summation is \[\sum_{n=1}^{40} -5+7(n-1)\]

OpenStudy (loser66):

Imitate the same process to get the answer for b) instead of +7, you do -10